The perimeter of a rectangle is 116centimeters. The length is .1meter graeter than twice the width. What is the length? show all your work.
Answer by rfer(16322)

2x+2(2x+10)=116
2x+4x+20=116
6x=96
x=16 cm
2x+10=42 cm